OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = x − 2, find f[g(−6)]. 9 3 −9 −3

hero (hero):

f(g(x)) = (x-2)-4 = x - 6 f(g(-6) = (-6-2)-4 = -8 - 4 = -12

hero (hero):

Answer choices are not correct

OpenStudy (anonymous):

i post 2 be a 1/2 betweetn (x) & x

hero (hero):

What?

OpenStudy (anonymous):

the fraction !/2 post 2 go between (x) & the = sign thats why its not comming out right

OpenStudy (anonymous):

1/2 i ment

hero (hero):

Well, fix your question and re-post it. That's all you have to do. Don't tell me what's wrong with your question, just simply fix it. There's an "Edit Question" button in the blue box area where you posted your question originally. Click that, then fix your question.

OpenStudy (anonymous):

it wont let me put the fraction ion the question thats why i was telling you where the fraction goes thats why its not comming out right tot he anwser choices i have.

hero (hero):

you don't know how to simply write (x-2)/2

hero (hero):

If that's what it is

OpenStudy (anonymous):

If f(x) = x − 4 and g(x) 1/2 = x − 2, find f[g(−6)].

hero (hero):

Okay, now that just doesn't make any sense

OpenStudy (anonymous):

the 1/2 part post 2 be a fraction

hero (hero):

Of course 1/2 is a fraction, but what you posted still makes no sense

hero (hero):

It's supposed to be f(x) = something and g(x) = something, not g(x) 1/2 = something

OpenStudy (anonymous):

oh yhea thats iht my fault